OpenStudy (anonymous):

The area of a rectangle is 133 square centimeters. If the perimeter is 52 centimeters, find the length and width of rectangle.

OpenStudy (anonymous):

xy = 133 --(1) 2(x+y)=52 --(2) from (1), x=133/y sub this into (2) 2(133/y +y)=52 133/y +y =26 133 + y^2 = 26y y^2 -26y -133=0 <-- solve by quadratic formula or calculator solve this and then put y into x=133/y to find x

OpenStudy (anonymous):

y = {30.378147197, -4.378147197}

OpenStudy (anonymous):

I don't think my answers right :/

OpenStudy (anonymous):

The length is 7 and the width is 19

OpenStudy (anonymous):

thank you so much!!!!